# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 7.2 Intertwining relations
The type I and the type II vertex operators satisfy the following relations.
$\Phi_{\nu}(\hbar^{1/2}v)L^{+}_{\mu\mu^{\prime\prime}}(u)=\sum_{\mu^{\prime}\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})_{\mu\nu}^{\mu^{\prime}\nu^{\prime}}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)\Phi_{\nu^{\prime}}(\hbar^{1/2}v),$ (7.7)
@@
LHS $=$ $\sum_{\mu^{\prime},\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)\chi(\hbar^{1/2}v/\tilde{v})^{-1}\Phi_{\nu^{\prime}}(\hbar^{1/2}v)\Psi^{*}_{\nu^{\prime\prime}}(\tilde{v}),{}$
RHS $=$ $\frac{\rho^{+}(u/v)}{\mu(u/v)}g\chi(\hbar^{1/2}u/\tilde{v})\chi(\hbar^{1/2}v/\tilde{v})^{-1}\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\nu}(\hbar^{1/2}v)\Psi^{*}_{\nu^{\prime\prime}}(\tilde{v})\Phi_{\mu}(\hbar^{1/2}u){}$
$=$ $\frac{\rho^{+}(u/v)}{\mu(u/v)}g\chi(\hbar^{1/2}u/\tilde{v})\chi(\hbar^{1/2}v/\tilde{v})^{-1}\chi(\hbar^{1/2}v/u)^{-1}\chi(\hbar^{1/2}u/\tilde{v})^{-1}{}$
×$\qquad\qquad\times\Phi_{\nu}(\hbar^{1/2}v)\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\mu}(\hbar^{1/2}u)\Psi^{*}_{\nu^{\prime\prime}}(\tilde{v}){}$
$=$ $\frac{\rho^{+}(u/v)}{\mu(u/v)}\chi(\hbar^{1/2}v/\tilde{v})^{-1}\chi(\hbar^{1/2}v/u)^{-1}\Phi_{\nu}(\hbar^{1/2}v)L^{+}_{\mu\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}(\tilde{v}).$
@@
LHS $=$ $\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\Phi_{\nu}(\hbar^{1/2}\tilde{v})\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})_{\mu^{\prime}\nu^{\prime}}^{\mu^{\prime\prime}\nu^{\prime\prime}},{}$
RHS $=$ $\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}g\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\Phi_{\nu}(\hbar^{1/2}\tilde{v})\Psi^{*}_{\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}({v})\Phi_{\mu}(\hbar^{1/2}u){}$
$=$ $\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}g\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\chi(\hbar^{1/2}u/v)^{-1}{}$
×$\qquad\qquad\times\Phi_{\nu}(\hbar^{1/2}\tilde{v})\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\mu}(\hbar^{1/2}u)\Psi^{*}_{\nu^{\prime\prime}}({v}){}$
$=$ $\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\chi(\hbar^{1/2}u/v)^{-1}\Phi_{\nu}(\hbar^{1/2}\tilde{v})L^{+}_{\mu\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}({v}).$
@@
$\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})_{\mu^{\prime}\nu^{\prime}}^{\mu^{\prime\prime}\nu^{\prime\prime}}{}$
$\quad=\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}u/v)^{-1}L^{+}_{\mu\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}({v}).$
@@
Finally let us consider a coalgebra structure which allows us to interpret the relations in Proposition 7.3 as intertwining relations on the modules of ${\cal U}=U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$. Let ${\cal L}^{+}(u)\in{\rm End}_{\mathbb{F}}({\cal F}^{(0,1)}_{u})\widetilde{\otimes}\;{\cal U}$ be the universal $L$-operator with
${\cal L}^{+}(u)\cdot{|\nu\rangle}_{u}\widetilde{\otimes}\eta=\sum_{\mu\in{\cal P}}{|\mu\rangle}_{u}\widetilde{\otimes}{\cal L}^{+}_{\mu\nu}(u)\eta,\qquad\eta\in{\cal U}.$
We assume ${\cal L}^{+}_{\mu\nu}(u)$ satisfies the $RLL$-relation (7.5). Let us further set ${\cal L}^{+}(u)={\cal L}^{+}(u;\mathfrak{z}^{*})e^{-c^{\perp}\widetilde{\otimes}Q}$ with
${\cal L}^{+}(u;\mathfrak{z}^{*})e^{-c^{\perp}\widetilde{\otimes}Q}\cdot{|\nu\rangle}_{u}\widetilde{\otimes}\eta=\sum_{\mu\in{\cal P}}{|\mu\rangle}_{u}\widetilde{\otimes}{\cal L}^{+}_{\mu\nu}(u;\mathfrak{z}^{*})e^{-Q}\eta.$
@@
$R^{+(12)}_{T^{1/2}}(u,v;\mathfrak{z}^{*}\hbar^{c^{\perp(3)}}){\cal L}^{+(1)}(u;\mathfrak{z}^{*}){\cal L}^{+(2)}(v;\mathfrak{z}^{*}\hbar^{c^{\perp(1)}}){}$
$\qquad\quad={\cal L}^{+(2)}(v;\mathfrak{z}^{*}){\cal L}^{+(1)}(u;\mathfrak{z}^{*}\hbar^{c^{\perp(2)}})R^{+*(12)}_{T^{1/2}}(u,v;\mathfrak{z}^{*}).$ (7.9)
We hence assume $e^{\pm Q}\cdot{|\lambda\rangle}_{u}={|\lambda\rangle}_{u}$ and
${\cal L}^{+}_{\mu\nu}(u)\cdot{|\lambda\rangle}_{v}={\cal L}^{+}_{\mu\nu}(u;\mathfrak{z}^{*})\cdot{|\lambda\rangle}_{v}=\sum_{\lambda^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z}^{*})_{\mu\lambda^{\prime}}^{\nu\lambda}{|\lambda^{\prime}\rangle}_{v},\qquad{|\lambda\rangle}_{v}\in{\cal F}^{(0,1)}_{v},$
${\cal L}^{+}_{\mu\nu}(u)\cdot\xi=L^{+}_{\mu\nu}(u)\xi,\qquad\xi\in{\cal F}^{(1,N)}_{\bullet}.$
@@
$\Delta({\cal L}^{+}_{\mu\nu}(u))=\sum_{\lambda}{\cal L}^{+}_{\mu\lambda}(u\gamma^{(2)})\widetilde{\otimes}{\cal L}^{+}_{\lambda\nu}(u),$ (7.10)
$\Delta(g(\mathfrak{z}))=g(\mathfrak{z})\widetilde{\otimes}1,\qquad\Delta(g(\mathfrak{z}^{*}))=1\widetilde{\otimes}g(\mathfrak{z}),\qquad\forall g(\mathfrak{z}),g(\mathfrak{z}^{*})\in{\mathbb{F}},$ (7.11)
$\Delta(\gamma^{1/2})=\gamma^{1/2}\widetilde{\otimes}\gamma^{1/2},\qquad\Delta(C)=C\widetilde{\otimes}C.$ (7.12)
We call this the standard comultiplication. Defining further a counit $\varepsilon$ and an antipode $S$ in terms of ${\cal L}^{+}(u)$ in the same way as in Sec.4.2 of [27], one can show that $(U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),\Delta,\varepsilon,\mu_{l},\mu_{r},S)$ is a Hopf algebroid.
@@
$\Phi(v){\cal L}^{+}_{\mu\nu}(u)=\Delta({\cal L}^{+}_{\mu\nu}(u))\Phi(v),$ (7.13)
${\cal L}^{+}_{\mu\nu}(u)\Psi^{*}(v)(\xi\widetilde{\otimes}{|\lambda\rangle}_{v})=\Psi^{*}(v)\Delta({\cal L}^{+}_{\mu\nu}(u))(\xi\widetilde{\otimes}{|\lambda\rangle}_{v}),\quad\xi\in{\cal F}^{(1,N)}_{\bullet},\ {|\lambda\rangle}_{v}\in{\cal F}^{(0,1)}_{v}.$ (7.14)
@@
$\mbox{RHS}\cdot\xi=\sum_{\mu^{\prime}}{\cal L}^{+}_{\mu\mu^{\prime}}(\gamma^{(2)}u)\widetilde{\otimes}{\cal L}^{+}_{\mu^{\prime}\nu}(u)\cdot\sum_{\lambda^{\prime}}{|\lambda^{\prime}\rangle}_{v}\widetilde{\otimes}\Phi_{\lambda^{\prime}}(v)\xi{}$
$\qquad=\sum_{\mu^{\prime},\lambda^{\prime},\lambda}R^{+}_{T^{1/2}}(\hbar^{1/2}u,v;\mathfrak{z}^{*})^{\mu^{\prime}\lambda^{\prime}}_{\mu\lambda}{|\lambda\rangle}_{v}\widetilde{\otimes}L^{+}_{\mu^{\prime}\nu}(u)\Phi_{\lambda^{\prime}}(v)\xi{}$
$\qquad=\sum_{\lambda}{|\lambda\rangle}_{v}\widetilde{\otimes}\sum_{\mu^{\prime},\lambda^{\prime}}R^{+}_{T^{1/2}}(\hbar^{1/2}u,v;\mathfrak{z})^{\mu^{\prime}\lambda^{\prime}}_{\mu\lambda}L^{+}_{\mu^{\prime}\nu}(u)\Phi_{\lambda^{\prime}}(v)\xi.$
@@
$\mbox{RHS}=\Psi^{*}(v)\cdot\sum_{\mu^{\prime}}{\cal L}^{+}_{\mu\mu^{\prime}}(\gamma^{(2)}u)\widetilde{\otimes}{\cal L}^{+}_{\mu^{\prime}\nu}(u)\cdot(\xi\widetilde{\otimes}{|\lambda\rangle}_{v}){}$
@@
$\qquad=\Psi^{*}(v)\cdot\sum_{\mu^{\prime},\lambda^{\prime}}R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*}\hbar^{-c^{\perp}(3)})^{\nu\lambda}_{\mu^{\prime}\lambda^{\prime}}L^{+}_{\mu\mu^{\prime}}(u)\xi\widetilde{\otimes}{|\lambda^{\prime}\rangle}_{v}{}$
$\qquad=\Psi^{*}(v)\cdot\sum_{\mu^{\prime},\lambda^{\prime}}L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*}{\hbar^{-c^{\perp(1)}-c^{\perp(3)}}})^{\nu\lambda}_{\mu^{\prime}\lambda^{\prime}}\xi\widetilde{\otimes}{|\lambda^{\prime}\rangle}_{v}{}$
@@
To obtain the fourth line, we used $\mathfrak{z}^{*}=\mathfrak{z}\hbar^{-c^{\perp}}$ and (2.15). Note also $p^{*}=p\hbar^{-1}$ on ${\cal F}^{(1,N)}_{\bullet}$. ∎

### v2

@@ 7.2. Intertwining relations
The type I and the type II dual vertex operators satisfy the following relations.
$\Phi_{\nu}(\hbar^{1/2}v)L^{+}_{\mu\mu^{\prime\prime}}(u)=\hskip-2.84544pt\sum_{\mu^{\prime}\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})_{\mu\nu}^{\mu^{\prime}\nu^{\prime}}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)\Phi_{\nu^{\prime}}(\hbar^{1/2}v),$ (7.5)
@@
[…]
$\hskip-28.45274pt\mbox{RHS}=\frac{\rho^{+}(u/v)}{\mu(u/v)}g\chi(\hbar^{1/2}u/\tilde{v})\chi(\hbar^{1/2}v/\tilde{v})^{-1}{}$
$\hskip 91.04872pt\times\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\nu}(\hbar^{1/2}v)\Psi^{*}_{\nu^{\prime\prime}}(\tilde{v})\Phi_{\mu}(\hbar^{1/2}u){}$
$\hskip-5.69046pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}g\chi(\hbar^{1/2}u/\tilde{v})\chi(\hbar^{1/2}v/\tilde{v})^{-1}\chi(\hbar^{1/2}v/u)^{-1}\chi(\hbar^{1/2}u/\tilde{v})^{-1}{}$
[…]
@@
$\hskip-14.22636pt\mbox{LHS}=\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\Phi_{\nu}(\hbar^{1/2}\tilde{v})\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})_{\mu^{\prime}\nu^{\prime}}^{\mu^{\prime\prime}\nu^{\prime\prime}},{}$
$\hskip-14.22636pt\mbox{RHS}=\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}g\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}\tilde{v}/v)^{-1}{}$
$\hskip 99.58464pt\times\Phi_{\nu}(\hbar^{1/2}\tilde{v})\Psi^{*}_{\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}({v})\Phi_{\mu}(\hbar^{1/2}u){}$
$\quad=\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}g\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\chi(\hbar^{1/2}u/v)^{-1}{}$
$\qquad\qquad\qquad\quad\qquad\times\Phi_{\nu}(\hbar^{1/2}\tilde{v})\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\mu}(\hbar^{1/2}u)\Psi^{*}_{\nu^{\prime\prime}}({v}){}$
$\hskip 8.5359pt=\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}\tilde{v}/v)^{-1}\chi(\hbar^{1/2}u/v)^{-1}{}$
$\hskip 99.58464pt\times\Phi_{\nu}(\hbar^{1/2}\tilde{v})L^{+}_{\mu\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}({v}).$
@@
$\hskip-14.22636pt\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})_{\mu^{\prime}\nu^{\prime}}^{\mu^{\prime\prime}\nu^{\prime\prime}}{}$
$=\frac{\rho^{+*}(u/v)}{\mu^{*}(u/v)}\chi(\hbar^{1/2}v/u)\chi(\hbar^{1/2}\tilde{v}/u)^{-1}\chi(\hbar^{1/2}u/v)^{-1}L^{+}_{\mu\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}({v}).$
@@
Finally let us consider a coalgebra structure which allows us to interpret the relations in Proposition 7.3 as intertwining relations on the modules of ${\mathcal{U}}=U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$. Let ${\mathcal{L}}^{+}(u)\in{\rm End}_{\mathbb{F}}({\mathcal{F}}^{(0,-1)}_{\bullet})\widetilde{\otimes}\;{\mathcal{U}}$ be the universal $L$-operator with
${\mathcal{L}}^{+}(u)\cdot{|\nu\rangle}_{u}\widetilde{\otimes}\eta=\sum_{\mu\in{\mathcal{P}}}{|\mu\rangle}_{u}\widetilde{\otimes}{\mathcal{L}}^{+}_{\mu\nu}(u)\eta,\qquad\eta\in{\mathcal{U}}.$
We assume ${\mathcal{L}}^{+}_{\mu\nu}(u)$ satisfies the $RLL$-relation in Proposition 7.2. Let us further set ${\mathcal{L}}^{+}(u)={\mathcal{L}}^{+}(u;\mathfrak{z}^{*})e^{c\widetilde{\otimes}Q}$ with
${\mathcal{L}}^{+}(u;\mathfrak{z}^{*})e^{c\widetilde{\otimes}Q}\cdot{|\nu\rangle}_{u}\widetilde{\otimes}\eta=\sum_{\mu\in{\mathcal{P}}}{|\mu\rangle}_{u}\widetilde{\otimes}{\mathcal{L}}^{+}_{\mu\nu}(u;\mathfrak{z}^{*})e^{-Q}\eta.$
@@
$R^{+(12)}_{T^{1/2}}(u,v;\mathfrak{z}^{*}\hbar^{-c^{(3)}}){\mathcal{L}}^{+(1)}(u;\mathfrak{z}^{*}){\mathcal{L}}^{+(2)}(v;\mathfrak{z}^{*}\hbar^{-c^{(1)}})$ (7.7)
$\qquad\quad={\mathcal{L}}^{+(2)}(v;\mathfrak{z}^{*}){\mathcal{L}}^{+(1)}(u;\mathfrak{z}^{*}\hbar^{-c^{(2)}})R^{+*(12)}_{T^{1/2}}(u,v;\mathfrak{z}^{*}).{}$
We hence assume $e^{\pm Q}\cdot{|\lambda\rangle}_{v}={|\lambda\rangle}_{v}$ for ${|\lambda\rangle}_{v}\in{\mathcal{F}}^{(0,-1)}_{v}$ and
${\mathcal{L}}^{+}_{\mu\nu}(u)\cdot{|\lambda\rangle}_{v}={\mathcal{L}}^{+}_{\mu\nu}(u;\mathfrak{z}^{*})\cdot{|\lambda\rangle}_{v}=\sum_{\lambda^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z}^{*})_{\mu\lambda^{\prime}}^{\nu\lambda}{|\lambda^{\prime}\rangle}_{v},{}$
${\mathcal{L}}^{+}_{\mu\nu}(u)\cdot\xi=L^{+}_{\mu\nu}(u)\xi,\qquad\xi\in{\mathcal{F}}^{(1,N)}_{\bullet}.$
@@
$\Delta({\mathcal{L}}^{+}_{\mu\nu}(u))=\sum_{\lambda}{\mathcal{L}}^{+}_{\mu\lambda}(u\gamma^{(2)})\widetilde{\otimes}{\mathcal{L}}^{+}_{\lambda\nu}(u),$ (7.8)
$\Delta(g(\mathfrak{z},p))=g(\mathfrak{z},p)\widetilde{\otimes}1,\quad\Delta(g(\mathfrak{z}^{*},p^{*}))=1\widetilde{\otimes}g(\mathfrak{z}^{*},p^{*}),\quad$ (7.9)
$\Delta(\gamma^{1/2})=\gamma^{1/2}\widetilde{\otimes}\gamma^{1/2},\qquad\Delta(C)=C\widetilde{\otimes}C$ (7.10)
for $\forall g(\mathfrak{z},p),g(\mathfrak{z}^{*},p^{*})\in{\mathbb{F}}$. We call this the standard comultiplication. Defining further a counit $\varepsilon$ and an antipode $S$ in terms of ${\mathcal{L}}^{+}(u)$ in the same way as in Sec.4.2 of [27], one can show that $(U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),\Delta,\varepsilon,\mu_{l},\mu_{r},S)$ is a Hopf algebroid.
@@
$\hskip-14.22636pt\Phi(v){\mathcal{L}}^{+}_{\mu\nu}(u)=\Delta({\mathcal{L}}^{+}_{\mu\nu}(u))\Phi(v),$ (7.11)
$\hskip-14.22636pt{\mathcal{L}}^{+}_{\mu\nu}(u)\Psi^{*}(v)(\xi\widetilde{\otimes}{|\lambda\rangle}_{v})=\Psi^{*}(v)\Delta({\mathcal{L}}^{+}_{\mu\nu}(u))(\xi\widetilde{\otimes}{|\lambda\rangle}_{v})$ (7.12)
for $\xi\in{\mathcal{F}}^{(1,N)}_{\bullet},\ {|\lambda\rangle}_{v}\in{\mathcal{F}}^{(0,-1)}_{v}$.
@@
$\mbox{RHS}\cdot\xi=\sum_{\mu^{\prime}}{\mathcal{L}}^{+}_{\mu\mu^{\prime}}(\gamma^{(2)}u)\widetilde{\otimes}{\mathcal{L}}^{+}_{\mu^{\prime}\nu}(u)\cdot\sum_{\lambda^{\prime}}{|\lambda^{\prime}\rangle}_{v}\widetilde{\otimes}\Phi_{\lambda^{\prime}}(v)\xi{}$
$\ \qquad\quad=\sum_{\mu^{\prime},\lambda^{\prime},\lambda}R^{+}_{T^{1/2}}(\hbar^{1/2}u,v;\mathfrak{z}^{*})^{\mu^{\prime}\lambda^{\prime}}_{\mu\lambda}{|\lambda\rangle}_{v}\widetilde{\otimes}L^{+}_{\mu^{\prime}\nu}(u)\Phi_{\lambda^{\prime}}(v)\xi{}$
$\ \qquad\quad=\sum_{\lambda}{|\lambda\rangle}_{v}\widetilde{\otimes}\sum_{\mu^{\prime},\lambda^{\prime}}R^{+}_{T^{1/2}}(\hbar^{1/2}u,v;\mathfrak{z})^{\mu^{\prime}\lambda^{\prime}}_{\mu\lambda}L^{+}_{\mu^{\prime}\nu}(u)\Phi_{\lambda^{\prime}}(v)\xi.$
@@
$\mbox{RHS}=\Psi^{*}(v)\cdot\sum_{\mu^{\prime}}{\mathcal{L}}^{+}_{\mu\mu^{\prime}}(\gamma^{(2)}u)\widetilde{\otimes}{\mathcal{L}}^{+}_{\mu^{\prime}\nu}(u)\cdot(\xi\widetilde{\otimes}{|\lambda\rangle}_{v}){}$
@@
$\ \qquad=\Psi^{*}(v)\cdot\sum_{\mu^{\prime},\lambda^{\prime}}R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*}\hbar^{c^{(3)}})^{\nu\lambda}_{\mu^{\prime}\lambda^{\prime}}L^{+}_{\mu\mu^{\prime}}(u)\xi\widetilde{\otimes}{|\lambda^{\prime}\rangle}_{v}{}$
$\ \qquad=\Psi^{*}(v)\cdot\sum_{\mu^{\prime},\lambda^{\prime}}L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*}{\hbar^{c^{(1)}+c^{(3)}}})^{\nu\lambda}_{\mu^{\prime}\lambda^{\prime}}\xi\widetilde{\otimes}{|\lambda^{\prime}\rangle}_{v}{}$
@@
To obtain the fourth line, we used $\mathfrak{z}^{*}=\mathfrak{z}\hbar^{c}$ and (2.6). Note also $p^{*}=p\hbar^{-1}$ on ${\mathcal{F}}^{(1,N)}_{\bullet}$. ∎
@@
It is a challenging problem to construct the universal ${\mathcal{L}}^{+}(u)$ satisfying the $RLL$ relation in Proposition 7.2 fully. One possible way along the idea discussed in this section is to construct the vertex operators both the type I and the type II dual on arbitrary level $(k,M)$ representation. For this purpose one needs to clarify the corresponding elliptic stable envelopes depending on $k$ as well as to find a realization of the $Z$-algebra part of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ studied in [32]. As in the elliptic quantum group associated with the affine Lie algebra $\widehat{\mathfrak{g}}$ [18, 27], we expect that ${\mathcal{L}}^{+}(u)$ and ${\mathcal{L}}^{-}(u)$ are not independent, namely one obtains the latter from the former by shifting the argument by $p^{*}\gamma\ (=p\gamma^{-1})$ essentially. Then one would recover $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ by considering the Gauss decomposition of them in the same way as Appendix C in [27].
It is also an interesting problem to consider the trigonometric limit $p\to 0$. In this limit $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$ tends to the quantum toroidal algebra $U_{t_{1},t_{2}}({\mathfrak{gl}}_{1,tor})$[36]. See [32] for detail. All the representations of the elliptic algebra have the definite trigonometric limits. In addition, the elliptic stable envelopes tend to those for the $\mathrm{K}$-theory of ${\mathcal{M}}(n,r)$ depending on the slope parameter, which appears as the certain limit of the Kähler parameter[1, 45]. Hence the trigonometric vertex operators and the $L^{+}$-operators depend on the slope. However it is not obvious to obtain the cohomological vertex functions from the $\mathrm{K}$-theoretic ones. One needs to take a scaling limit of the equivariant parameters $t_{1},t_{2},u$ further.